OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). A. y = -4x + 1 B. y = -1/2 x + 9 C. y = 4x + 1 D. y = -2x + 9

OpenStudy (anonymous):

its D

OpenStudy (the_fizicx99):

\[M = \frac{ (Y2-Y1) }{ (X2-X1) }\] \[ (x,y) \rightarrow (4,1)\] \[(x,y) \rightarrow (5,-1)\] \[M = \frac{ -1-1 }{5-4} \rightarrow \frac{ -2 }{1} \] M = -2/1 Now that you have the slope, use the formula: \[y-y1=M(x-x1)\] to write the slope intercept equation. \[Y-1=\frac{ -2 }{ 1 }(x-4)\] \[y - 1= \frac{ -2 }{ 1}x + 8\] \[y= \frac{ -2 }{ 1 }x + 9\]